# HashAttention: Semantic Sparsity for Faster Inference

### v1

@@ 7 Conclusion
HashAttention proposes a principled approach to sparsify attention based on hashing key-value pairs and queries to a compressed semantic space. Near neighbour comparisons can be efficiently performed in this semantic space using bit operations. The quality of tokens retreived by HashAttention, under similar resources, is significantly superior to other sparsity baselines, leading to upto 4×4$\times$4 × further reduction of tokens as compared to leading sparse attention baselines. This leads to 4×4$\times$4 × gains in efficiency over other baselines.
Appendix A Design Choices
In this section, we discuss some of the design choices for our method.
A.1 Retrieval Algorithm
One way to look at HashAttention is that, given a query it divides the set of keys into subsets according to levels of quantized similarities. For instance, if we use 32 bit signatures, then we only have access to 32 levels of similarity. Under such a similarity assignment, what is the best way to choose retrieved set of keys? There are two natural approaches
  1. 1.
    depth-based-retrieval(d): We find the maximum bit match score, say $d_{max}$, and retrieve all the keys with scores $\geq\lambda_{d}=d_{max}-d$
    $d_{max}=\max_{k\in K}\{\textrm{bit-match}(q,k)\}$ (9)
    $\lambda_{d}=d_{max}-d$ (10)
  2. 2.
    num-based-retrieval(n): We find the bit match score of the $n^{th}$ (max) key, say $\lambda_{n}$. Then all the keys with scores $\geq\lambda_{n}$ are chosen.
  3. 3.
    combined(n,d): we use the maximum of the two scores to avoid retrieving unnecessary excess tokens.
    $\lambda_{d,n}=\max(\lambda_{d},\lambda_{n})$ (11)
A.2 Note on #retrieved tokens, scaling, variability and such
An important question in designing the retrieval is whether, we should always output a fixed number of tokens or have some variability of tokens. Semantics dictate that depending on the query, the number of tokens important for attention mechanism can be different, and thus, we allow the design to retrieve variable number of tokens for every query.
We observe that using only depth-based-retrieval leads to linear scaling of count of retrieved tokens. (i.e. number of retrieved tokens increases linearly with context size). Using only number based retrieval causes unnecessary retrieval in some cases when only a few tokens have high bit match score – especially true in small context lengths. Thus we use combination of both as our retrieval algorithm. We see a sublinear scale up with the combined recipe.
Appendix B Evaluation Philosophy
How do we evaluate sparse attention? Two schools of thought exist depending on what we are trying to achieve. For end to end trained models, ofcourse sparse attention needs to be applied at all times as defacto thing to do. On the other hand, when trying to speed up decoding for existing LLM models, it maybe considered that prefilling is done offline and thus sparse attention only needs to be applied to decoding phase. However, most baselines which follow the second line of thought for evaluation often disregard that we might need to interact with this long term context differently – i.e. we might want to ask different questions on the same document. In this case, it would be ideal to apply full attention on long term context ( e.g. document) but apply sparse attention to local prompt (e.g. question) and following decoding. This is especially important in evaluation using datasets. For instance, consider passage retrieval task. In this case, you only need to predict the paragraph number. So if you apply full attention to the entire prompt, the comparison of question and retrieval from context has already happened with full attention and thus it does not capture the quality of attention.
To remedy the situation agnostically, we offset the sparse attention application – meaning we start applying sparse attention to the end of the prompt before entering decode phase. In our experiments we use a offset of 128 tokens.
Appendix C Baselines (Adapted-HashAttention)
@@ C.6 Double Sparsity
Appendix D Best Sparse solution for adaptation
Using the following notations, $\mathbf{V}:n\times d$ value matrix, $\mathbf{a}:n\times 1$ attention scores. $\mathbf{S}:n\times n$ diagonal sampling matrix. Then under the sampling the final embedding is,

### v2

@@ 7 Conclusion
HashAttention proposes a principled approach to sparsify attention based on hashing key-value pairs and queries to a compressed semantic space. Near neighbor comparisons can be efficiently performed in this semantic space using bit operations. The quality of tokens retrieved by HashAttention, under similar resources, is significantly superior to other sparsity baselines, leading to upto 4×4$\times$4 × further reduction of tokens as compared to leading sparse attention baselines. Overall, we see reduction in KV Cache usage upto 1632×16-32$\times$16 - 32 × with minimal drop in quality (within $1$ point) across benchmarks and models, leading to latency and throughput gains.
Acknowledgments
Sky Computing Lab is supported by gifts from Accenture, AMD, Anyscale, Cisco, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Lightspeed, Mibura, Microsoft, NVIDIA, Samsung SDS, and SAP. Aditya Desai’s position at sky-lab is supported by Mayfield Charitable.
Impact Statement
This paper presents work whose goal is to advance the efficiency of Machine Learning which can lead to wider adoption of large language models. There are many potential societal consequences of advancing machine learning and its efficiency, none of which we feel must be specifically highlighted here.
Appendix A Baselines
@@ A.6 Double Sparsity
Appendix B Lemma Proofs
B.1 Lemma 1: Best Sparse solution for adaptation for 1-token selection
Using the following notations, $\mathbf{V}:n\times d$ value matrix, $\mathbf{a}:n\times 1$ attention scores. $\mathbf{S}:n\times n$ diagonal sampling matrix with exactly one non-zero in the diagonal (selection of a single token). Note that this ignores the cross token corrleation in value vectors. Then under the sampling the final embedding is,
@@ B.1 Lemma 1: Best Sparse solution for adaptation for 1-token selection
B.2 Lemma 2: Identifying topk w.r.t isolated token score is a MIPS problem
$a_{i}||v_{i}||_{2}$ $=\frac{e^{\langle\mathbf{q},\mathbf{k}\rangle}||v_{i}||_{2}}{\sum e^{\langle\mathbf{q},\mathbf{k}\rangle}}$ (11)
$=\frac{e^{\langle\mathbf{q},\mathbf{k}\rangle+\log(||\mathbf{v}_{i}||)}}{\sum e^{\langle\mathbf{q},\mathbf{k}\rangle}}$ (12)
$\propto\langle\mathbf{q},\mathbf{k}\rangle+\log(||\mathbf{v}_{i}||)$ (13)
$=\langle[\mathbf{q},1],[\mathbf{k},\log(||\mathbf{v}_{i}||)]$ (14)
Appendix C Other ablations
C.1 Recall rates of HashAttention and DS
C.2 Detailed pareto curves for each dataset averaged in the main paper.
The pareto curves are shown in Figure 5
Figure 5: Pareto curves at 32 bits PTPA for HashAttention, DS and Quest
Table 4: Longbench (non-chinese tasks)
[…]
HashAttention-16x 53.65 100 42.82 43.52 29.16 34.99 26.68 70.29 51.92 1.85 24.75 42.97 92.14 27.93 64.89 60.49 48.00 51.08
Full 54.68 100 43.75 45.80 29.41 34.90 27.11 71.43 55.40 8.29 25.14 45.62 90.55 30.19 62.87 55.31 48.78 51.48
Table 5: Longbench (chinese tasks). The hash attention is trained on english data only
multifieldqa_zh dureader vcsum passage-retrieval-zh lsht AVG
59.31 30.93 17.48 91.71 41.71 48.228
61.36 34.51 17.14 96.86 44.57 50.888
C.3 Latency of SCORE computation
Table 6: In order to understand the number of bits upto which HashAttention would show latency gains, we compare the full inner product computation latency with bitwise score computation latency for varying number of bits We can see that upto 512 bits, our bit-wise kernel is superior to full inner product computation.
Latency of SCORE computation
Inner product Hamming distance using packed 64 bit integers
#tokens 128 x 16Float 64bit 128bit 256bit 512bit
262144 0.13 0.085 0.084 0.087 0.112
524288 0.266 0.085 0.086 0.117 0.218
1048576 0.619 0.087 0.129 0.226 0.429
2097152 1.698 0.144 0.247 0.445 0.85
4194304 3.272 0.28 0.485 0.88 1.694
8388608 6.733 0.552 0.96 1.755 3.378
16777216 8.306 1.1 1.911 3.504 6.746
33554432 16.61 2.188 3.817 7.005 13.48
The latency comparison of SCORE computation and full inner product computation is presented in Table 6
C.4 Does HashAttention resolve the out-of-distribution issue of queries and keys
While HashAttention does not directly aim to resolve the out-of-distribution, the learned mappings do cause the distributions to be closer. We present some data in Table 7
Table 7: Average Cosine similarity between query and key embeddings of top-32 tokens. This is computed for Llama-3.1-8B-Instruct model
Average Cosine Similarity of top-32 tokens
Layer number original embeddings HashAttention (tanh) HashAttention(sign)
0 -0.111818 0.1777 0.180198
4 -0.119951 0.4922 0.179186
8 -0.109657 0.2676 0.1802
12 -0.10329 0.1309 0.178136
16 -0.110954 0.2422 0.182292
20 -0.0985931 0.2188 0.179659
24 -0.114441 0.293 0.179346
28 -0.101307 0.2441 0.179191
C.5 Learned bit representations performed in HashAttention vs. data-agnostic bit signatures using LSH (random signed projections) (Gionis et al., 1999)
Table 8: If we use LSH (used in MagicPig(Chen et al., 2024)) instead of learned bit signatures via mapping functions as done in HashAttention, the bit signature sizes needed for LSH are considerably higher. This is expected since LSH does not exploit the data distribution in the selection of projections
Compression #bits passage_retrieval_en multifieldqa_en hotpotqa triviaqa Average
16x HashAttention 32 99.43 55.18 52.34 93.32 75.0675
16x LSH 32 48.57 32.62 34.38 80.02 48.8975
LSH 256 89.14 47.97 53.97 88.07 69.7875
LSH 512 85.14 50.38 51.42 87.6 68.635
LSH 1024 91.43 49.75 55.25 OOM
8x LSH 32 69.14 42.71 46.63 81.78 60.065
LSH 256 82.86 50.18 52.06 90.3 68.85
LSH 512 95.43 52.64 54.85 87.51 72.6075
LSH 1024 98.86 51.84 54.67 OOM
If we use LSH instead of learned bit signatures via mapping functions as done in HashAttention, the bit signature sizes needed for LSH are considerably higher. This is expected since LSH does not exploit the data distribution in the selection of projections. The results on LongBenchmark datasets is presented in Table 8. This is consistent with results from MagicPig(Chen et al., 2024).
C.6 Quality vs. bits
The Table 9 shows the cross entropy loss (lower the better) for top-64 prediction while using different bit-widths.
Table 9: Cross-entropy loss for different embedding dimensions evaluated on 250 × 32K samples.
Dimension Cross Entropy Loss @ 250 × 32K samples
16 0.243
32 0.212
64 0.204